# MiniMaxAD: A Lightweight Autoencoder for Feature-Rich Anomaly Detection

### v3

@@ Abstract
Previous unsupervised anomaly detection (UAD) methods often struggle to handle the extensive diversity in training sets, particularly when they contain stylistically diverse and feature-rich samples, which we categorize as feature-rich anomaly detection datasets (FRADs). This challenge is evident in applications such as multi-view and multi-class scenarios. To address this challenge, we developed MiniMaxAD, a lightweight autoencoder designed to efficiently compress and memorize extensive information from normal images. Our model employs a technique that enhances feature diversity, thereby increasing the effective capacity of the network. It also utilizes large kernel convolution to extract highly abstract patterns, which contribute to efficient and compact feature embedding. Moreover, we introduce an Adaptive Contraction Hard Mining Loss (ADCLoss), specifically tailored to FRADs. In our methodology, any dataset can be unified under the framework of feature-rich anomaly detection, in a way that the benefits far outweigh the drawbacks. Our approach has achieved state-of-the-art performance in multiple challenging benchmarks.
@@ 1 Introduction
Anomaly detection presents a significant challenge in the field of computer vision. Owing to the difficulties associated with collecting and labeling anomalous samples, unsupervised learning methods have become increasingly favored for anomaly detection. These methods utilize only normal samples for training and are engineered to identify anomalous images and segment anomalous regions during inference. Recently, UAD technology is increasingly used in more complex scenarios to meet growing demand.
Previous anomaly detection methods often require training a separate model for each class of objects, which can lead to significant memory consumption. In response, UniAD (You et al. 2022) proposes a unified setting that trains a single model on multiple classes of normal samples. They also highlight the limitations of the traditional one-category-one-model approach, especially in cases where normal samples exhibit considerable intra-class diversity. This challenge is especially pronounced in the fields of unmanned supermarkets and multi-view anomaly detection, as illustrated by the newly introduced GoodsAD(Zhang et al. 2024) and Real-IAD(Wang et al. 2024) datasets, which exhibit significant intra-class diversity.
In this study, we revisited the existing UAD dataset and categorized it based on the diversity of samples within the training set, a concept we will refer to as intra-set diversity. We define datasets with substantial intra-set diversity as Feature-Rich Anomaly Detection Datasets (FRADs), in contrast to Feature-Poor Anomaly Detection Datasets (FPADs), as illustrated in Figure 1. Notably, FPADs can be easily converted to FRADs using the unified setting. This diversity necessitates that neural networks extensively memorize normal features. Among FRADs, PatchCore (Roth et al. 2021) demonstrates superior performance in applications such as unmanned supermarkets (Zhang et al. 2024) and multi-view scenarios (Wang et al. 2024), benefiting from an efficient memory bank that stores normal contexts. However, this approach also significantly increases inference times due to its reliance on time-consuming query operations. Additionally, as the number of training samples grows, so does the memory bank’s size, complicating the application of PatchCore in unified settings. Consequently, we aim to develop a high-performance, lightweight UAD method that addresses the challenges of FRADs effectively.
@@
Methods based on multi-scale reconstruction (Deng and Li 2022; Guo et al. 2023; Tien et al. 2023; Zhang et al. 2023b) are celebrated for their simplicity and efficiency. These methods use a paired encoder and decoder to reconstruct features, effectively capturing the manifold of normal samples. During testing, areas with higher reconstruction errors are flagged as anomalies. However, they face two primary challenges: (a) the network’s limited information capacity may inadequately reconstruct normal areas, resulting in false positives; (b) its robust generalization ability might allow it to accurately reconstruct abnormal areas, leading to false negatives, a phenomenon often referred to as the “identical shortcut” (You et al. 2022). Experimental results from GoodsAD (Zhang et al. 2024) show that these methods tend to generate a high rate of false positives by mistakenly classifying many normal foreground regions as abnormal in FRADs. Moreover, replacing the backbone of Reverse Distillation (RD) (Deng and Li 2022) with a larger model, such as WideResNet-101 (Zagoruyko and Komodakis 2016), significantly reduces performance, as illustrated in Figure 2(a). This suggests that dual challenges of limited capacity and the “identical shortcut” are present when dealing with FRADs.
In response to the dual challenges, our study introduces a core motivation: maximizing the “effective capacity” of a lightweight autoencoder. Our approach is built on an efficient and compact architecture, the Reverse Distillation (RD) Architecture. However, it suffers from severe performance degradation when confronted with FRADs, particularly on the GoodsAD dataset, where the decline is critical—it nearly fails to reconstruct any foreground regions of the objects. To address this, we adopted a more aggressive improvement strategy by starting from its fundamental architecture. Our initial step is to reduce the number of parameters, which offers the dual benefits of enhancing computational efficiency and mitigating the “identical shortcut” issue. Although reducing the number of parameters typically results in decreased information capacity due to fewer neurons, we recognize that the ratio of information capacity to the number of parameters may have been underestimated. Despite having fewer parameters, a significant number of features can still be retained. To achieve this, we utilize large kernel convolutions to extract more comprehensive representations with relatively fewer parameters. Additionally, we employ a technique to prevent neuron inactivation, GRN (Woo et al. 2023), which enhances network capacity by reactivating saturated or inactive neurons, thereby enabling them to represent more information. Finally, we introduce a novel Adaptive Contraction Hard Mining Loss (ADCLoss). Unlike previous hard mining methods, our approach allows the mining rate to adapt dynamically based on the characteristics of the given dataset. Our main contributions are summarized below:
@@
  • We developed a novel Adaptive Contraction Hard Mining Loss (ADCLoss) tailored for FRADs with excellent performance.
@@ 2.1 Unsupervised Anomaly Detection
Methods based on reconstruction. In unsupervised anomaly detection (UAD), methods based on image-level reconstruction train models to reconstruct normal images, leveraging the assumption that anomalies will result in poor reconstruction quality due to model unfamiliarity with normal images. Recently, OCR-GAN (Liang et al. 2023) improved the performances by decoupling input images into various frequency components, employing multiple generators for enhanced reconstruction . On the feature level, the introduction of the teacher-student model (Bergmann et al. 2020; Gu et al. 2023; Deng and Li 2022; Batzner, Heckler, and König 2023; Guo et al. 2023), inspired by knowledge distillation (Hinton, Vinyals, and Dean 2015), utilizes a pre-trained network as a ’teacher’ to guide a ’student’ model trained only on normal images. This approach targets feature reconstruction. Recently, MambaAD(He et al. 2024a), based on reconstruction, explored the application of Mamba(Gu and Dao 2023) to unsupervised anomaly detection (UAD).
Methods Based on K-Nearest Neighbor (KNN). KNN-based methods leverage feature matching for inference, as demonstrated in various studies (Cohen and Hoshen 2020; Nazare, de Mello, and Ponti 2018; Napoletano, Piccoli, and Schettini 2018; Cohen, Tzachor, and Hoshen 2023). One prominent example, PatchCore (Roth et al. 2021), utilizes a large pre-trained network to extract normal features and applies a core-set downsampling technique to mitigate storage and inference costs. This approach has exhibited strong performance and robustness across diverse industrial anomaly detection datasets. Despite these advantages, the costs associated with inference rise incrementally as the training sample size increases, due to the expanding memory bank.
@@ 2.2 Hard Mining Strategies
Hard mining strategies are widely employed in reconstruction methods to enhance model performance. The core idea behind these strategies is to focus the model’s optimization efforts on the regions of the image that are difficult to reconstructt, typically the foreground regions. IKD (Cao et al. 2022) introduces a hard mining method that adaptively adjusts the margin based on the statistical information (mean and variance) of feature distances within each mini-batch. This method uses a fixed discard rate as a hyperparameter, discarding feature points with smaller reconstruction errors. ReContras (Guo et al. 2023) refines this approach by making the discard rate variable, decreasing it over the course of iterations to achieve a coarse-to-fine optimization process. Instead of discarding the feature points themselves, ReContras discards the gradients associated with those points. EfficientAD (Xiao et al. 2023) takes a more aggressive approach by discarding 99.9% of the feature points in the feature map, focusing exclusively on fine-grained optimization.
@@ 3.1 Reverse Distillation for Anomaly Detection
Reverse Distillation (RD) (Deng and Li 2022) is an important method in UAD, featuring a streamlined architecture that consists of a frozen pre-training encoder (teacher), a bottleneck, and a feature reconstruction decoder (student). The encoder extracts multi-scale spatial features from the input image, which the bottleneck then condenses to minimize redundancy. The decoder, designed as an inverse counterpart to the encoder, reconstructs the encoder’s output to achieve feature-level alignment. When faced with abnormal samples during testing, the decoder’s inability to accurately reconstruct the encoder’s output serves to highlight discrepancies, thereby generating an anomaly map. Formally, let $f_{E}^{k},f_{D}^{k}\in\mathbb{R}^{C_{k}\times H_{k}\times W_{k}}$ represent the feature maps output by the encoder and decoder at the $k$-th level, respectively. Here, $C_{k}$, $H_{k}$, and $W_{k}$ denote the number of channels, the height, and the width of the output feature map at each level. The training objective is to minimize the global cosine distance loss between the corresponding output features of each layer for normal images, quantified by the following formula:
@@ 3.2 Expanding the Effective Capacity Cap
The Global Response Normalization (GRN) unit can reactivate saturated or dead neurons, thereby increasing neuron utilization and enhancing model capacity. Specifically, the GRN unit boosts channel contrast and selectivity by aggregating features at the channel level. This is achieved using the L2-norm on the feature map $X_{i}$, followed by normalization to compute the feature normalization coefficient $n_{x}$:
@@ 3.3 MinMax Thinking
Our core motivation can be addressed in two ways: (1) to mitigate the “identical shortcut” phenomenon by reducing the number of parameters, and (2) to maximize the effective capacity of a lightweight autoencoder under the constraints imposed by (1). Here, “effective” highlights the model’s ability to accurately reconstruct normal foreground regions, while “capacity” refers to the model’s enhanced information density. Based on this premise, an intuitive strategy is to extract deeper, more abstract representations to achieve highly compact information storage. Consequently, large kernel convolution emerges as a powerful method for implementing this strategy. In recent visual recognition research, advances in large kernel convolution have renewed interest in convolutional networks. UniRepLKNet (Ding et al. 2023) proposes a series of design guidelines for large-kernel convolutions; they have developed a powerful backbone equipped with a GRN Unit that perfectly fits our methodology. This network comprises several downsampling blocks and two distinct types of blocks—the Large Kernel Block (LarK Block) and the Small Kernel Block (SmaK Block). As shown in the encoder section of the Figure 3(a). Specifically, $E^{1}$ uses a 3$\times$3 convolution kernel in the SmaK Block, while $E^{2}$ and $E^{3}$ utilize a 13$\times$13 convolution kernel in the LarK Block. The full architecture of the proposed model is depicted in Figure 3(a).
To illustrate the substantial advantages of MiniMaxAD over RD in handling FRADs, we conducted a quantitative analysis of the feature map alignment process, where the Decoder reconstructs the Encoder’s features. We perform the L2-norm on the feature maps of the encoder and decoder, after which the variance was calculated as $\frac{1}{3}\sum_{k=1}^{3}\sigma({f^{k}_{E}/\left\|f^{k}_{E}\right\|})^{2}$ and $\frac{1}{3}\sum_{k=1}^{3}\sigma({f^{k}_{D}/\left\|f^{k}_{D}\right\|})^{2}$. As depicted in Figure 4(a), benefiting from the GRN unit and large kernel design, the variance of MiniMaxAD is about five times that of RD, indicating that MiniMaxAD can characterize richer information. We conduct a more analysis of this process from the perspective of entropy, detailed in the Appendix.
@@ 3.4 Adaptive Contraction Strategy
4 Experiments
4.1 Experimental Settings
Datasets. MVTec AD and VisA feature high-resolution images across multiple object and texture categories, highlighting both surface and structural defects. AeBAD-S is specifically focused on domain diversity within aero-engine blade images. GoodsAD captures extensive intra-class diversity with 6,124 high-resolution images of common supermarket goods, showcasing a single product offered by multiple brands. Real-IAD is a recent, large UAD dataset comprising 150K images across 30 categories, with each category containing five default perspectives. Detailed information is presented in Appendix.
Implementation. Network optimization was carried out using an AdamW optimizer (Loshchilov and Hutter 2019), with the $\beta$ parameters set to (0.9, 0.999) and a weight decay of 0.00005. The learning rate was set at 0.005, accompanied by an ExponentialLR scheduler with a gamma of 0.995. The batch size for all experiments was 16. For experiments in the single-class setting, GoodsAD was trained for 240 epochs, ViSA for 150 epochs, AeBAD-S for 130 epochs, and Real-IAD for 50 epochs. In the multi-class setting, MVTec AD and ViSA were trained for 160 epochs, while Real-IAD was trained for only 1 epoch. Results were reported from the last epoch for each scenario. All experiments were performed on an NVIDIA RTX 4090 GPU using PyTorch 2.0. The model does not use efficient large kernel convolution as suggested by (Ding et al. 2023).
Metrics. We use two of the most commonly employed metrics for anomaly detection and segmentation as our basic evaluation metrics(Bergmann et al. 2019): image-level Area Under the Receiver Operating Characteristic Curve (I-AUROC) and pixel-level Area Under the Per-Region Overlap (AUPRO). Additionally, following previous work(Zhang et al. 2023a; He et al. 2024a), we adopt supplementary metrics on certain datasets for a more comprehensive evaluation, including Average Precision (AP) (Zavrtanik, Kristan, and Skocaj 2021) and F1-score-max (F1_max) (Zou et al. 2022). We also calculate the average of the seven evaluation metrics mentioned above, denoted as mAD, to represent the model’s overall performance.
4.2 Anomaly Detection and Segmentation on FPADs
@@
PCore RD EfficientAD Ours-fp
I-AUROC 95.1 96.0 98.1 96.1
AUPRO 91.2 70.9 94.0 90.4
@@
In this section, we focus on demonstrating the performance of MiniMaxAD on FPADs using the two most commonly used metrics (see Section 4.1). We compare our method with RD(Deng and Li 2022), PatchCore(Ishida et al. 2023), MMR(Zhang et al. 2023b), and EfficientAD(Batzner, Heckler, and König 2023), all of which are powerful state-of-the-art (SoTA) methods.

### v4

@@ Abstract
Previous industrial anomaly detection methods often struggle to handle the extensive diversity in training sets, particularly when they contain stylistically diverse and feature-rich samples, which we categorize as feature-rich anomaly detection datasets (FRADs). This challenge is evident in applications such as multi-view and multi-class scenarios. To address this challenge, we developed MiniMaxAD, a efficient autoencoder designed to efficiently compress and memorize extensive information from normal images. Our model employs a technique that enhances feature diversity, thereby increasing the effective capacity of the network. It also utilizes large kernel convolution to extract highly abstract patterns, which contribute to efficient and compact feature embedding. Moreover, we introduce an Adaptive Contraction Hard Mining Loss (ADCLoss), specifically tailored to FRADs. In our methodology, any dataset can be unified under the framework of feature-rich anomaly detection, in a way that the benefits far outweigh the drawbacks. Our approach has achieved state-of-the-art performance in multiple challenging benchmarks. Code is available at: https://github.com/WangFengJiee/MiniMaxAD
@@ 1 Introduction
Anomaly detection plays a crucial role in quality control within modern intelligent manufacturing and has attracted growing attention in recent years. Due to the challenges in collecting and labeling anomaly samples, a growing number of researchers are reducing their reliance on supervisory signals [4]. As a result, unsupervised learning methods are becoming increasingly popular for anomaly detection. These methods utilize only normal samples for training and are engineered to identify anomalous images and segment anomalous regions during inference. Recently, unsupervised anomaly detection (UAD) technology is increasingly used in more complex scenarios to meet growing demand.
Previous anomaly detection methods often require training a separate model for each class of objects, which can lead to significant memory consumption. In response, UniAD [37] proposes a unified setting that trains a single model on multiple classes of normal samples. They also highlight the limitations of the traditional one-category-one-model approach, especially in cases where normal samples exhibit considerable intra-class diversity. This challenge is especially pronounced in the fields of unmanned supermarkets and multi-view anomaly detection, as illustrated by the newly introduced GoodsAD [41] and Real-IAD [31] datasets, which exhibit significant intra-class diversity.
In this study, we revisited the existing UAD dataset and categorized it based on the diversity of samples within the training set, a concept we will refer to as intra-set diversity. We define datasets with substantial intra-set diversity as Feature-Rich Anomaly Detection Datasets (FRADs), in contrast to Feature-Poor Anomaly Detection Datasets (FPADs), as illustrated in Figure 1. Notably, FPADs can be easily converted to FRADs using the multi-class setting. This diversity necessitates that neural networks extensively memorize normal features. Among FRADs, PatchCore [27] demonstrates superior performance in applications such as unmanned supermarkets [41] and multi-view scenarios [31], benefiting from an efficient memory bank that stores normal contexts. However, this approach also significantly increases inference times due to its reliance on time-consuming query operations. Additionally, as the number of training samples grows, so does the memory bank’s size, complicating the application of PatchCore in unified settings. Consequently, we aim to develop a high-performance, effective UAD method that addresses the challenges of FRADs effectively.
@@
Methods based on multi-scale reconstruction [8, 13, 30, 43] are celebrated for their simplicity and efficiency. These methods use a paired encoder and decoder to reconstruct features, effectively capturing the manifold of normal samples. During testing, areas with higher reconstruction errors are flagged as anomalies. However, they face two primary challenges: (a) the network’s limited information capacity may inadequately reconstruct normal areas, resulting in false positives; (b) its robust generalization ability might allow it to accurately reconstruct abnormal areas, leading to false negatives, a phenomenon often referred to as the “identical shortcut” [37]. Experimental results from GoodsAD [41] show that these methods tend to generate a high rate of false positives by mistakenly classifying many normal foreground regions as abnormal in FRADs. Moreover, replacing the backbone of Reverse Distillation (RD) [8] with a larger model, such as WideResNet-101 [38], significantly reduces performance, as illustrated in Figure 2(a). This suggests that dual challenges of limited capacity and the “identical shortcut” are present when dealing with FRADs.
In response to the dual challenges, our study introduces a core motivation: maximizing the “effective capacity” of a relatively lightweight autoencoder. Our approach is built on an efficient and compact architecture, the Reverse Distillation (RD) Architecture. However, it suffers from severe performance degradation when confronted with FRADs, particularly on the GoodsAD dataset, where the decline is critical—it nearly fails to reconstruct any foreground regions of the objects. To address this, we adopted a more aggressive improvement strategy by starting from its fundamental architecture. Our initial step is to reduce the number of parameters, which offers the dual benefits of enhancing computational efficiency and mitigating the “identical shortcut” issue. Although reducing the number of parameters typically results in decreased information capacity due to fewer neurons, we recognize that the ratio of information capacity to the number of parameters may have been underestimated. Despite having fewer parameters, a significant number of features can still be retained. To achieve this, we utilize large kernel convolutions to extract more comprehensive representations with relatively fewer parameters. Additionally, we employ a technique to prevent neuron inactivation, GRN [33], which enhances network capacity by reactivating saturated or inactive neurons, thereby enabling them to represent more information. Finally, we introduce a novel Adaptive Contraction Hard Mining Loss (ADCLoss). Unlike previous hard mining methods, our approach allows the mining rate to adapt dynamically based on the characteristics of the given dataset. Our main contributions are summarized below:
@@
  • We developed a novel ADCLoss tailored for FRADs with excellent performance.
@@ 2.1 Unsupervised Anomaly Detection
Methods based on reconstruction. In unsupervised anomaly detection (UAD), methods based on image-level reconstruction train models to reconstruct normal images, leveraging the assumption that anomalies will result in poor reconstruction quality due to model unfamiliarity with normal images. Recently, OCR-GAN [21] improved the performances by decoupling input images into various frequency components, employing multiple generators for enhanced reconstruction. TDAD [32] employs advanced diffusion models to achieve high-quality reconstruction. On the feature level, the introduction of the teacher-student model [3, 12, 8, 1, 13], inspired by knowledge distillation [17], utilizes a pre-trained network as a ’teacher’ to guide a ’student’ model trained only on normal images. This approach targets feature reconstruction. Recently, MambaAD [15], based on reconstruction, explored the application of Mamba [11] to unsupervised anomaly detection (UAD).
Methods based on simulated anomalies. This approach, which simulates anomalies to provide supervisory signals, includes various innovative methods [30, 22, 42, 35, 28]. For instance, [20] utilizes a technique where a portion of an image is cut and randomly pasted elsewhere to create forged anomalies. Moreover, DRAEM [39] employs Perlin noise to construct anomalies and segments these during training to generalize the segmentation capability to actual anomalies. More recently, MemSeg [36] attempted to simulate anomalies within the foreground region of images. However, the segmentation of the foreground region proved to be unstable. Furthermore, SimpleNet [22] introduces gaussian noise into the feature space to simulate anomalies. The effectiveness of these methods hinges on the extent to which the pseudo anomalies resemble actual anomalies and on the placement of these fabricated anomalies. Generalizing these methods to new datasets remains challenging, and there is considerable scope for improvement.
Methods Based on K-Nearest Neighbor (KNN). KNN-based methods leverage feature matching for inference, as demonstrated in various studies [6, 26, 25, 7]. One prominent example, PatchCore [27], utilizes a large pre-trained network to extract normal features and applies a core-set downsampling technique to mitigate storage and inference costs. This approach has exhibited strong performance and robustness across diverse industrial anomaly detection datasets. Despite these advantages, the costs associated with inference rise incrementally as the training sample size increases, due to the expanding memory bank.
@@ 2.2 Hard Mining Strategies
Hard mining strategies are widely employed in reconstruction methods to enhance model performance. The core idea behind these strategies is to focus the model’s optimization efforts on the regions of the image that are difficult to reconstructt, typically the foreground regions. IKD [5] introduces a hard mining method that adaptively adjusts the margin based on the statistical information (mean and variance) of feature distances within each mini-batch. This method uses a fixed discard rate as a hyperparameter, discarding feature points with smaller reconstruction errors. ReContrast [13] refines this approach by making the discard rate variable, decreasing it over the course of iterations to achieve a coarse-to-fine optimization process. Instead of discarding the feature points themselves, ReContras discards the gradients associated with those points. EfficientAD [34] takes a more aggressive approach by discarding 99.9% of the feature points in the feature map, focusing exclusively on fine-grained optimization.
@@ 3.1 Reverse Distillation for Anomaly Detection
Reverse Distillation (RD) [8] is an important method in UAD, featuring a streamlined architecture that consists of a frozen pre-training encoder (teacher), a bottleneck, and a feature reconstruction decoder (student). The encoder extracts multi-scale spatial features from the input image, which the bottleneck then condenses to minimize redundancy. The decoder, designed as an inverse counterpart to the encoder, reconstructs the encoder’s output to achieve feature-level alignment. When faced with abnormal samples during testing, the decoder’s inability to accurately reconstruct the encoder’s output serves to highlight discrepancies, thereby generating an anomaly map. Formally, let $f_{E}^{k},f_{D}^{k}\in\mathbb{R}^{C_{k}\times H_{k}\times W_{k}}$ represent the feature maps output by the encoder and decoder at the $k$-th level, respectively. Here, $C_{k}$, $H_{k}$, and $W_{k}$ denote the number of channels, the height, and the width of the output feature map at each level. The training objective is to minimize the global cosine distance loss between the corresponding output features of each layer for normal images, quantified by the following formula:
@@ 3.2 Expanding the Effective Capacity Cap
The Global Response Normalization (GRN) [33] can reactivate saturated or dead neurons, thereby increasing neuron utilization and enhancing model capacity. Specifically, the GRN unit boosts channel contrast and selectivity by aggregating features at the channel level. This is achieved using the L2-norm on the feature map $X_{i}$, followed by normalization to compute the feature normalization coefficient $n_{x}$:
@@ 3.3 MinMax Thinking
Our core motivation can be addressed in two ways: (1) to mitigate the “identical shortcut” phenomenon by reducing the number of parameters, and (2) to maximize the effective capacity of a lightweight autoencoder under the constraints imposed by (1). Here, “effective” highlights the model’s ability to accurately reconstruct normal foreground regions, while “capacity” refers to the model’s enhanced information density. Based on this premise, an intuitive strategy is to extract deeper, more abstract representations to achieve highly compact information storage. Consequently, large kernel convolution emerges as a powerful method for implementing this strategy. In recent visual recognition research, advances in large kernel convolution have renewed interest in convolutional networks. UniRepLKNet [10] proposes a series of design guidelines for large-kernel convolutions; they have developed a powerful backbone equipped with a GRN Unit that perfectly fits our methodology. This network comprises several downsampling blocks and two distinct types of blocks—the Large Kernel Block (LarK Block) and the Small Kernel Block (SmaK Block). As shown in the encoder section of the Figure 3(a). Specifically, $E^{1}$ uses a 3$\times$3 convolution kernel in the SmaK Block, while $E^{2}$ and $E^{3}$ utilize a 13$\times$13 convolution kernel in the LarK Block. The full architecture of the proposed model is depicted in Figure 3(a).
To illustrate the substantial advantages of MiniMaxAD over RD in handling FRADs, we conducted a quantitative analysis of the feature map alignment process, where the Decoder reconstructs the Encoder’s features. We perform the L2-norm on the feature maps of the encoder and decoder, after which the variance was calculated as $\frac{1}{3}\sum_{k=1}^{3}\sigma({f^{k}_{E}/\left\|f^{k}_{E}\right\|})^{2}$ and $\frac{1}{3}\sum_{k=1}^{3}\sigma({f^{k}_{D}/\left\|f^{k}_{D}\right\|})^{2}$. As depicted in Figure 4(a), benefiting from the GRN unit and large kernel design, the variance of MiniMaxAD is about five times that of RD, indicating that MiniMaxAD can characterize richer information. We conduct a more analysis of this process from the perspective of entropy, detailed in the Appendix. Additionally, Figure 2(b) shows that while RD’s global cosine distance loss is lower than MiniMaxAD’s, RD’s Image-level AUROC is significantly worse. This suggests that the reduced feature diversity in WideResNet-50 simplifies the decoder’s alignment task, whereas MiniMaxAD’s greater feature diversity better supports complex anomaly detection.
@@ 3.4 Adaptive Contraction Strategy
PCore RD EAD MambaAD Ours-fp
I-AUROC 95.1 96.0 98.1 94.7 96.1
AUPRO 91.2 70.9 94.0 91.6 90.4
@@
4 Experiments